# Rosette formations as symmetry-breaking events: theory and experiment

### v1

@@ II.3 Nematic interactions boosts the probability of observing rosettes
Next, we investigated the effect of nematic interactions, modeled via the addition of a term to the equation dictating the orientation of the self-propulsion $\theta_{i}$ that enforces the alignment between the direction of the self-propulsion velocity and the principal axis of elongation of the cell:
$\dot{\theta}_{i}=\eta_{i}-J\sin(\theta_{i}-\phi_{i})$ (5)
@@
The angular dynamics is controlled by the interplay of rotational diffusion at rate $D_{r}$ and alignment at rate $J$, whose inverse $\tau_{J}=\frac{1}{J}$ is the response time required by the cell to reorient its polarization in the direction of the resultant force exerted by its neighbors. Figure 3 shows stationary snapshots of the resulting simulations. In particular, Figure 3a displays representative stationary configurations of populations of cells having different motility in presence of an external harmonic potential of intensity $k=0.2$ with ($J>0)$ and without ($J=0$) nematic interaction. While increasing the motility tends to impede the formation of rosette, alignment allows high motile cells to organize in rosette-like assemblies. In addition, increasing the strength of the alignment interaction boosts the probability of observing transient rosette at low values of target shape factor (see Figure 3b).
@@
Figure 6: Measuring dynamical features of neural precursor cells. a) Bright field (left) and RFP fluorescence (right) microscopy images of iPSCs seeded a low density. Yellow and cyan arrows mark the direction and modulus of the cell self velocity and polarization, respectively. Green line shows the trajectory of the highlighted cell during the time lapse. b) Cell displacement as a function of the difference between displacement’ and polarization’s angles. c) Mean squared displacement as a function of time for a representative cell dynamics. A sketch of the extracted quantities is reported together with the best fit of Eq. 7. d) Box plot representation of the distribution of measured self velocity (cyan) and persistence times (red).
@@
Figure 7: Neural rosette formation process. a) Confocal fluorescence microscopy image of iPSCs expressing the FUS protein fused to a RFP. Red lines represent the best Voronoi tassellation considering cell nuclei as seeds. b) Comparison between snapshot taken during the temporal evolution of a iPSCs population and a dynamcal Voronoi simulation in presence of an harmonic potential. Transient rosettes are highlighted in gray. c) Average shape factor as a function of time for cell population that form or do not form rosette structures. Representative snapshots of the considered regions are shown on top and bottom of the panel for a rosette-forming region and a no-rosette region, respectively. d) Confocal fluorescence microscopy image of a lumen-centered rosette observed after 11 days from the starting of the differentiation procedure on the iPSC cells. Nuclear RFP FUS fluorescence is represented in red, while N-cadherin is in cyan. Scale bar 20 $\mu m$.
@@
II.4 Alignment increases the size of lumen-centered rosettes and influence its polarization
Transient rosettes are known to either resolve or stabilize in lumen centered polarized rosettes Harding et al. (2014). Rosette lumens are central cavities formed during epithelial morphogenesis. They arise when a group of polarized epithelial cells organizes in a circular (rosette) pattern, a process critical for organ development, including in the kidneys, mammary glands, and neural tube Martin-Belmonte and Mostov (2008); Bryant and et al. (2010). The epithelial cells surrounding the lumen are highly polarized, possessing distinct apical and basal domains. The apical membrane is enriched with markers such as aPKC, Par3, and tight junction proteins like ZO-1, which are essential for establishing and maintaining polarity Martin-Belmonte and Mostov (2008). The lumen is initially filled with extracellular fluid and secreted proteins (e.g., mucins and glycoproteins). In some cases, apoptotic cells are cleared from the center, contributing to the formation of the luminal space Bryant and et al. (2010). The basement membrane underlying the epithelial cells is rich in extracellular matrix proteins, including laminin, collagen IV, nidogen, and perlecan. This matrix provides structural support and transduces biochemical signals that regulate cell polarity and proliferation O’Brien and et al. (2001). Cytoskeletal components such as actin filaments, microtubules, and myosin II are critical in regulating cell shape and polarity. Additionally, cell-cell adhesion is maintained through tight junctions (e.g., ZO-1), adherens junctions (e.g., E-cadherin), and desmosomes, which together stabilize the rosette structure Martin-Belmonte and Mostov (2008). Multiple signaling pathways are involved in lumen formation. Pathways such as Wnt, Notch, Hedgehog, and BMP provide spatial cues, while regulators like Rho GTPases and the Par complex directly influence cell polarity and lumen expansion Bryant and et al. (2010). We modeled the net effect of the lumen on nearby cells using a Lennard-Jones potential irradiating from one of the Voronoi cell centers. In particular, we imposed that one of the N cells composing the population has zero velocity, vanishingly small target area and perimeter and exerts a potential energy interaction with the other cells of the population of the form:
@@ II.4 Alignment increases the size of lumen-centered rosettes and influence its polarization
Figure 4a displays a snapshot of a stationary configuration obtained via the Active Voronoi model in presence of the Lennard-Jones potential with parameters, $\epsilon_{0}=1.0$ and $\sigma_{0}=3$, where cells forming the first, second and third layer around the central lumen are colored in green, gold and blue, respectively. Qualitatively, it can be seen that cells organize into a polarized structure around the lumen. Notably, this behavior is conserved upon varying the lumen size (Figure 4b). From a more quantitative point of view, we characterized the size and organization of the rosette by counting the number of cells that compose the various layers and their relative orientation with respect to the lumen center. In particular, we defined the n-th layer size as $l_{n}=\sum_{i}\delta(d_{i0}-n)$, where $d_{i0}$ represents the minimum number of cells one as to go through to reach the center of the Lennard-Jones potential starting from cell $i$. Moreover, we introduced a polarization index for the n-th layer of cells, $\phi_{n}$, as $\phi_{n}=\braket{\Delta_{i}}_{n}$, where $\Delta_{i}$ is defined as the angle between the elongation angle of the cell and the direction of the vector linking the i-th cell center and the center of the lumen. Figure 4c shows the number of cells in the first cellular layers, $l_{1}$ as a function of time for different values of the cell target shape factor, $p_{0}$. After an initial transient time, the number of cells in the layer reaches a stationary state, whose value depends on the target shape factor of the cell population: the higher the target shape factor, the higher the number of cells in the layer. This behavior is evident looking at the stationary values of $l_{1}$ in Figure 4d as a function the cell target shape factor, $p_{0}$ and Lennard-Jones effective radius, $\sigma_{0}$ for two values of the Lennard-Jones potential strength, $\epsilon_{LJ}$. The number of cells in the rosette increases not only changing the size of the lumen, as one would expect, but also the strength of the Lennard-Jones potential, i.e., the attraction of the cells toward the lumen. The increase of cells around the lumen reflects in a polarization of the layer with respect to the lumen center (see Figure 4e,g). As one can see from the snapshots in Figure 4f, the size and polarization of the rosette increases according to the values of the cell target shape factor, $p_{0}$, for a given lumen size. Notably, the number of cells in a given rosette layer and the corresponding polarization of the layer with respect to the lumen are correlated. Intuitively, the more cells in a layer, the more they will tend to assume an elongated form. This trend is appreciable by comparing the behavior of the average polarizations of cells in the first cellular layers in Figure 4g-h with those in panels c) and d) of the same Figure. The presence of elongated cells suggests that self-alignment interaction may influence the dynamics of the rosette. To probe this hypothesis, we repeat simulations turning on the nematic interaction. As one can see both by visual inspection of the configurations in panel i) of Figure 4 and by comparing the distributions of $l_{n}$ and $\phi_{n}$ in Figure 4l (for a system of 400 cells having a target shape factor of 3.9), the presence of self-alignment interaction significantly increases the number of cells in the first and second layers of the rosette. Indeed, these layers are composed by $\sim 30$ more cells than their correspectives obtained for $J=0$. This difference reduces to $\sim 10$ for the third layer.
@@ II.5 Leader cell-centered rosettes display similar behaviors without a potential term
So far, we modeled the presence/effect of the lumen through an effective potential energy term, under the assumption that the concentration of actin filaments in the lumen creates an effective steric hindrance of the nearby cells and a net attraction, with cells attempting to reach the lumen region. To test to what extend our findings are dictated by the presence of the potential, we carried our simulations in with potential is turned off while the leader cells have different target areas, $A_{0}$. Figure 5a displays snapshots of stationary configurations obtained via the Active Voronoi model in the presence of a leader cell of target area $A_{0}$. Top row displays snapshots for a population of cells having target shape factor $p_{0}=3.2$, while the bottom row corresponds to cells with $p_{0}=3.8$. Qualitatively, configurations are similar to those observed in the presence of the LJ potential, especially for target areas higher than the target areas of the cells composing the population. Notably, the introduction of a leader cell with a bigger target area provokes the formation of transient rosettes in the close proximity of the leader-cell centered rosette. The presence of such a cell introduces a defect in the system, which mimics the perturbation produced by the external potential: to accommodate the defect, nearby cells tend to form transient rosette structures.
Evaluating the stationary number of cells and the polarization index for the first cellular layer, we found that rosettes formed around a leader cell manifests a behavior qualitatively similar to those formed in presence of a LJ potential. In particular, looking at $l_{1}$ (green shaded map) and $\phi_{1}$ (blue shaded map) as a function the cell target shape factor, $p_{0}$ and leader cell target area, $A_{0}$, we see that both the size and the polarization of the rosette increases as the population passes from solid-like to liquid-like as observed for the LJ case (see Figure 5b-c) Finally, we probed the effect of self-alignment, retrieving a similar, albeit weak, effect of the this interaction on both the size and structure of the rosette. Results are shown in Figure 5d-e for the rosette first layers in absence and presence of self-alignment interaction for different values of cells target shape factor, $p_{0}$.
@@ II.6 Neural rosettes are observed in fluidity-preserving 2D cultures
To test our model predictions, we collected fluorescence microscopy time-lapses during the neural differentiation of iPSCs (see Methods for details). Under the proper differentiation conditions, iPSCs form rosettes after 8-10 days Karus et al. (2014). In the first 4-6 days from the plating, cells are in a low-density conditions, thus its is possible to directly measure the dynamical behavior of single cells. Figure 6a shows a snapshot of the dynamics of iPSCs tagged with a fluorescent marker (RFP) fused to a nuclear enriched protein (FUS). Note that fluorescence of the nuclei allows for an accurate tracking of the cellular motion, while low density eases the segmentation of all cells in the field of view. Analyzing the dynamics of single cells, we found that they preferentially move along the direction of their elongation as can be seen by looking at Figure 6b, that displays the instantaneous displacement as a function of the angle between the motion direction and the elongation axis. Next, we evaluated the mean squared displacement (MSD) as function of time for all identified trajectories (see Methods). Figure 6c shows an example of experimental MSD together with the best fit of the expected theoretical behavior trend:
$MSD(t)=2v_{0}^{2}\left(\frac{t}{\tau}-1+e^{-\frac{t}{\tau}}\right)$ (7)
@@
For this trajectory, the change from balistic to diffusive motion is evident from the modification of the slopes at $\tau\sim 250$ s. Results for all tracked cells are reported as box plots in Figure 6d. Cells display a self-propulsion of $v_{0}=0.4\pm 0.1$ $\mu m/s$ and a persistence time of $\tau=260\pm 50$ s. The resulting persistence length is $l_{p}=v_{0}\tau\sim 100\mu m$. In order to compare the behavior of the real data with respect to their in silico counterparts, we proceeded to determine the distance scaling factor as follows: we took representative snapshots of the confluent cell population cropping images in LxL square regions (see Figure 7a), compute the corresponding Voronoi tesselation (see Methods) and determined the ratio, $f$, between the effective edge length of the region (measured in $\mu m$) and the squared root of the number of cells present in the region ($\sqrt{N_{c}}$). Note that simulations were performed considering N cells in a square box of $\sqrt{N}$ edge length. This allows us to properly rescale the measured quantities and compare them with simulations. Rescaling of the simulation box equals to have cells with an effective length of $\sim\sqrt{A_{0}}=1$, while the effective length in the data is $l_{c}\sim 10\mu m$. Thus, $\tilde{l}_{p}=f\cdot l_{p}=10l_{c}$.
@@
Having fixed the persistence times of the rotational dynamics to $\tau=50$, simulations can be compared to real data for values of the self propulsion velociry around $0.2$.
@@
Indeed, analyzing the tesselation more in details, we found that several snapshots displays cells in rosette-like arrangements, qualitatively similar to those observed in the simulations performed with the harmonic potential. Figure 7b shows some examples of real (top row) and simulated (bottom row) configurations. In particular, it is interesting to note that such transient rosettes are not isolated but form clusters as we observed in the simulations with cell being in the fluid-like region. Measuring the shape factors of the cells in the snapshots, we found an average shape factor always higher than 3.95, indicating that those configurations are in a fluid like regime even with their relatively small self propulsion velocities. Notably, as shown in Figure 7c, we found that cells in regions that do not form rosettes tend to reduce their shape factor in time, while regions on the layer that maintain a shape factor above 3.95 display rosettes that mature in lumen centered structures (Figure 7d). This trend is in accordance with our numerical predictions: the probability of observing transient rosettes is higher for cells in the fluid-like regions, where also the rosette size and polarization is maximum in the lumen centered case (see Figure 4).

### v2

@@ II.3 Self-alignment interactions boost the probability of observing rosettes
Next, we investigated the effect of self-alignment interactions, modeled via the addition of a term to the equation dictating the orientation of the self-propulsion $\theta_{i}$, which enforces the alignment between the direction of the self-propulsion velocity and the principal axis of elongation of the cell:
$\dot{\theta}_{i}=-J\sin(\theta_{i}-\phi_{i})+\eta_{i}$ (5)
@@
The angular dynamics is controlled by the interplay of rotational diffusion at rate $D_{r}$ and alignment at rate $J$, whose inverse $\tau_{J}=\frac{1}{J}$ is the response time required by the cell to reorient its polarization in the direction of the resultant force exerted by its neighbors Paoluzzi et al. (2021). Figure 3 shows stationary snapshots of the resulting simulations. In particular, Figure 3a displays representative stationary configurations of populations of cells having different motility in the presence of an external harmonic potential of intensity $k=0.2$ with ($J>0)$ and without ($J=0$) self-alignment interaction. While increasing the motility tends to impede the formation of rosettes, alignment allows even highly motile cells to organize in rosette-like assemblies. In addition, increasing the strength of the alignment interaction boosts the probability of observing transient rosettes at low values of target shape factor (see Figure 3b).
@@
Figure 6: Measuring dynamical features of neural precursor cells. a) Bright field (left) and RFP fluorescence (right) microscopy images of iPSCs seeded at low density. Yellow and cyan arrows mark the direction and modulus of the cell self-velocity and polarization, respectively. The green line shows the trajectory of the highlighted cell during the time lapse. b) Cell displacement as a function of the difference between the displacement and polarization angles. c) Mean squared displacement as a function of time for a representative cell dynamics. A sketch of the extracted quantities is reported together with the best fit of Eq. 7. d) Box plot representation of the distribution of measured self velocity (cyan) and persistence times (red).
@@
Figure 7: Neural rosette formation process. a) Confocal fluorescence microscopy image of iPSCs expressing the FUS protein fused to a RFP. Red lines represent the best Voronoi tassellation considering cell nuclei as seeds. b) Comparison between snapshots taken during the temporal evolution of an iPSCs population and a dynamic Voronoi simulation in the presence of a harmonic potential. Transient rosettes are highlighted in gray. c) Average shape factor as a function of time for cell populations that form or do not form rosette structures. Representative snapshots of the considered regions are shown at the top and bottom of the panel for a rosette-forming region and a no-rosette region, respectively. d) Confocal fluorescence microscopy image of a lumen-centered rosette observed after 11 days from the start of the differentiation procedure on the iPSC cells. Nuclear RFP FUS fluorescence is represented in red, while N-cadherin is in cyan. Scale bar 20 $\mu m$.
@@
II.4 Alignment increases the size of lumen-centered rosettes and influences their polarization
Transient rosettes can mature in lumen-centered ones Harding et al. (2014). We modeled the net effect of the lumen on nearby cells using a Lennard-Jones potential irradiating from one of the Voronoi cell centers. In particular, we imposed that one of the N cells composing the population has zero velocity, vanishingly small target area and perimeter, and exerts a potential energy interaction with the other cells of the population of the form:
@@ II.4 Alignment increases the size of lumen-centered rosettes and influences their polarization
Figure 4a displays a snapshot of a stationary configuration obtained via the Active Voronoi model in the presence of the LJ potential with parameters, $\epsilon_{0}=1.0$ and $\sigma_{0}=3$. Cells forming the first, second, and third layers around the central lumen are colored in green, gold, and blue, respectively. Qualitatively, it can be seen that cells organize into a polarized structure around the lumen. Notably, this behavior is conserved upon varying the lumen size (Figure 4b). To get quantitative insight into the role of the model parameters, we characterized the size and organization of the rosette by counting the number of cells that compose the various layers and their relative orientation with respect to the lumen center. In particular, we defined the n-th layer size as $l_{n}=\sum_{i}\delta(d_{i0}-n)$, where $d_{i0}$ represents the minimum number of cells one has to go through to reach the center of the Lennard-Jones potential starting from cell $i$. Moreover, we introduced a polarization index for the n-th layer of cells, $\phi_{n}$, as $\phi_{n}=\braket{\Delta_{i}}_{n}$, where $\Delta_{i}$ is defined as the angle between the elongation angle of the cell and the direction of the vector linking the i-th cell center and the center of the lumen. Figure 4c shows the number of cells in the first cellular layers, $l_{1}$, as a function of time for different values of the cell target shape factor, $p_{0}$. After an initial transient time, the number of cells in the layer reaches a stationary state, whose value depends on the target shape factor of the cell population: the higher the target shape factor, the higher the number of cells in the layer. This behavior is evident looking at the stationary values of $l_{1}$ in Figure 4d as a function of the cell target shape factor, $p_{0}$, and Lennard-Jones effective radius, $\sigma_{0}$, for two values of the Lennard-Jones potential strength, $\epsilon_{LJ}$. The number of cells in the rosette increases, not only changing the size of the lumen, as one would expect, but also the strength of the Lennard-Jones potential, i.e., the attraction of the cells toward the lumen. The increase of cells around the lumen reflects in a polarization of the layer with respect to the lumen center (see Figure 4e,g). As one can see from the snapshots in Figure 4f, the size and polarization of the rosette increase according to the values of the cell target shape factor, $p_{0}$, for a given lumen size. Notably, the number of cells in a given rosette layer and the corresponding polarization of the layer with respect to the lumen are correlated. Intuitively, the more cells in a layer, the more they will tend to assume an elongated form. This trend is appreciable by comparing the behavior of the average polarizations of cells in the first cellular layers in Figure 4g-h with those in panels c) and d) of the same Figure. The presence of elongated cells suggests that self-alignment interaction may influence the dynamics of the rosette. To probe this hypothesis, we repeat simulations turning on the self-alignment interaction. As one can see both by visual inspection of the configurations in panel i) of Figure 4 and by comparing the distributions of $l_{n}$ and $\phi_{n}$ in Figure 4l (for a system of 400 cells having a target shape factor of 3.9), the presence of self-alignment interaction significantly increases the number of cells in the first and second layers of the rosette. Indeed, these layers are composed of $\sim 30$ % more cells than their counterparts obtained for $J=0$. This difference reduces to $\sim 10$ % for the third layer.
@@ II.5 Leader cell-centered rosettes display similar behaviors without a potential term
So far, we modeled the presence/effect of the lumen through an effective potential energy term, under the assumption that the concentration of actin filaments in the lumen creates an effective steric hindrance of the nearby cells and a net attraction, with cells attempting to reach the lumen region. To test to what extent our findings are dictated by the presence of the potential, we carried out simulations in which the potential is turned off while a leader cell has a different target area, $A^{L}_{0}$, with respect to the rest of the population. Figure 5a displays snapshots of stationary configurations obtained via the Active Voronoi model in the presence of a leader cell of target area $A^{L}_{0}$. The top row displays snapshots for a population of cells having target shape factor $p_{0}=3.2$, while the bottom row corresponds to cells with $p_{0}=3.8$. Qualitatively, configurations are similar to those observed in the presence of the LJ potential, especially for target areas higher than the target areas of the cells composing the population. Notably, the introduction of a leader cell with a bigger target area provokes the formation of transient rosettes in the close proximity of the leader-cell-centered rosette. The presence of such a cell introduces a defect in the system, which mimics the perturbation produced by the external potential: to accommodate the defect, nearby cells tend to form transient rosette structures.
Evaluating the stationary number of cells and the polarization index for the first cellular layer, we found that rosettes formed around a leader cell manifest a behavior qualitatively similar to those formed in the presence of a LJ potential. In particular, looking at $l_{1}$ (green shaded map) and $\phi_{1}$ (blue shaded map) as a function the cell target shape factor, $p_{0}$ and leader cell target area, $A_{0}$, we see that both the size and the polarization of the rosette increases as the population passes from solid-like to liquid-like as observed for the LJ case (see Figure 5b-c) Finally, we probed the effect of self-alignment, retrieving a similar, albeit weak, effect of this interaction on both the size and structure of the rosette. Results are shown in Figure 5d-e for the rosette first layers in the absence and presence of self-alignment interaction for different values of cell target shape factor, $p_{0}$.
@@ II.6 Neural rosettes are observed in fluidity-preserving 2D cultures
To test our model predictions, we collected fluorescence microscopy time-lapses during the neural differentiation of iPSCs (see Methods for details). Under the proper differentiation conditions, iPSCs form rosettes after 8-10 days Karus et al. (2014). In the first 4-6 days from the plating, cells are in low-density conditions, thus it is possible to directly measure the dynamical behavior of single cells. Figure 6a shows a snapshot of the dynamics of iPSCs tagged with a fluorescent marker (RFP) fused to a nuclear-enriched protein (FUS). Note that fluorescence of the nuclei allows for an accurate tracking of the cellular motion, while low density eases the segmentation of all cells in the field of view (see Methods). Analyzing the dynamics of single cells, we found that they preferentially move along the direction of their elongation. This can be appreciated by looking at Figure 6b, which displays the instantaneous displacement as a function of the angle between the motion direction and the elongation axis. Next, we evaluated the mean squared displacement $\langle\Delta r^{2}\rangle$ as a function of time for all identified trajectories (see Methods). Figure 6c shows an example of experimental $\langle\Delta r^{2}\rangle$ together with the best fit of the expected theoretical behavior $\langle\Delta r^{2}\rangle_{th}$ in the case of active particles moving with self-propulsion velocity $v_{0}$ and characterized by a persistence time $\tau$ Bechinger et al. (2016)
$\langle\Delta r^{2}\rangle_{th}=2v_{0}^{2}\left(\frac{t}{\tau}-1+e^{-\frac{t}{\tau}}\right)\;.$ (7)
@@
According to (7), the cell performs a persistent motion characterized by a ballistic regime in the $\langle\Delta r^{2}\rangle$ on short time scales, i.e., for $t\ll\tau$. In the ballistic regime, $\langle\Delta r^{2}\rangle$ grows quadratically $\langle\Delta r^{2}\rangle\propto v_{0}^{2}t^{2}$. At larger time scales, i. e., $\tau\gg\tau$, the random walk is characterized by a diffusive behavior so that $\langle\Delta r^{2}\rangle$ grows linearly $\langle\Delta r^{2}\rangle\propto D\,t$, with $D=v_{0}^{2}\tau/2$ in the case of active motion in two spatial dimensions. As it is easy to check, these two regimes can be obtained by considering a small and large time expansion of (7). $\tau$ sets the crossover time between the two regimes. For this trajectory, the change from ballistic to diffusive motion is evident from the modification of the slopes at $\tau\sim 250$ s. The results of all tracked cells are reported as box plots in Figure 6d. Cells show a self-propulsion of $v_{0}=0.4\pm 0.1$ $\mu m/s$ and a persistence time of $\tau=260\pm 50$ s. The resulting persistence length is $l_{p}=v_{0}\tau\sim 100\mu m$. In order to compare the behavior of the real data with respect to their in silico counterparts, we proceeded to determine the distance scaling factor as follows: we took representative snapshots of the confluent cell population cropping images in LxL square regions (see Figure 7a), compute the corresponding Voronoi tesselation (see Methods) and determined the ratio, $f$, between the effective edge length of the region (measured in $\mu m$) and the squared root of the number of cells present in the region ($\sqrt{N_{c}}$). Note that simulations were performed considering N cells in a square box of $\sqrt{N}$ edge length. This allows us to properly rescale the measured quantities and compare them with simulations. Rescaling of the simulation box equals having cells with an effective length of $\sim\sqrt{A_{0}}=1$, while the effective length in the data is $l_{c}\sim 10\mu m$. Thus, $\tilde{l}_{p}=f\cdot l_{p}=10l_{c}$.
@@
Having fixed the persistence times of the rotational dynamics to $\tau=50$, simulations can be compared to real data for values of the self-propulsion velocity around $0.2$.
@@
Indeed, analyzing the tesselation more in detail, we found that several snapshots display cells in rosette-like arrangements, qualitatively similar to those observed in the simulations performed with the harmonic potential. Figure 7b shows some examples of real (top row) and simulated (bottom row) configurations. In particular, it is interesting to note that such transient rosettes are not isolated but form clusters, as we observed in the simulations with cells being in the fluid-like region. Measuring the shape factors of the cells in the snapshots, we found an average shape factor always higher than 3.95, indicating that those configurations are in a fluid-like regime. Notably, as shown in Figure 7c, we found that cells in regions that do not form rosettes tend to reduce their shape factor in time, while regions on the layer that maintain a shape factor above 3.95 display rosettes that mature in lumen-centered structures (Figure 7d). This trend is in accordance with our numerical predictions: the probability of observing transient rosettes is higher for cells in the fluid-like regions, where also the rosette size and polarization are maximum in the lumen-centered case (see Figure 4).